SLAC T-510 News
This page lists news from the experiment.
14 Feb 2014 - Status as of Feb 14
- Took data with thin target.
- Finished up near field data taking.
- Took additional ICT data with the transformer closer to the target and with collimators at 10 at 100 mm.
- Might not get the beam tomorrow.
The peak response of the cumulative sum of the ICT as a function of input charge. The linearity is remarkable. Another way to think about this value, is the integral of the ICT response curve up to the zero crossing.
13 Feb 2014 - Status as of Feb 13
UHF off-axis measurements and near field data is taken.
12 Feb 2014 - Status as of Feb 12
- Beam is on around 10am
- Running tests to adjust the scope settings for ICT, scintillator, S-band horn.
- Took a lot of ICT calibration data.
- UHF vertical scan of the Cherenkov cone completed
11 Feb 2014 - Status as of Feb 8-11
- Working on DAQ tuning and ICT setup and calibration.
- Took some B-field measurements.
- UHF antenna tower is at the Cherenkov angle.
- Building search is copmleted. Ready for the beam.
28 Jan 2014 - Status as of Jan 28
ESA is in permitted access. Target is surveyed and radiation boundary is removed. Post beam calibration procedures:
- B-field linearity check
- cable calibration
- some DAQ tweaking to increase speed and stability in preparation for the February run.
- pulsing antenna array at the Cherenkov cone
- review and survey the experiment's geometry
B-field on the beam axis. Current is divided among three sections of coils.
27 Jan 2014 - Status as of Jan 27
- Beam is at 11 am. Starting VHF scan.
- Strong CW at 164 MHz. Source unknown.
- Opened a collimator to increase the current. CW does not bother anymore.
- Beam stopped at 9 pm sharp.
- Finished VHF scan. Done taking data.
27 Jan 2014 - Status as of Jan 26
Success! With good beam all day scanned the vertically across the Cherenkov cone. V-pol and H-hol track the cone. Eh-pol/Ev-pol vs B-field is linear. Data looks good, follows theoretical predictions. We have an experiment. Congratulations to everyone!
27 Jan 2014 - Status as of Jan 25
In between the beam tune up for LCLS and SLAC hardware problems, only got about 1 hour of the beam. Resolved issues with flaky scope channels. Test runs 12 through 36. Will live with occasional sync slips. Turned on B-field. Can reach 840G. See strong H-pol signal!
24 Jan 2014 - Status as of Jan 24
- Cabled all scope channels to the antennas
- Fixed some DAQ hiccups
- Beam is intermittent, waiting for it to come back
Diagram of the experiment
23 Jan 2014 - Status as of Jan 23
- More DAQ debugging from the counting house
- ESA closed at 9:55 am for the search
- Search completed at 1:00 pm
- Badges for people on site added to website
- Ready for the first beam shots
- After a long wait due to radiological survey the beam is finally on
- See a strong signal in V-pol (see visualized data in Channel 3). H-pol is not yet connected
Mapped magnetic field
22 Jan 2014 - Status as of Jan 22
- More calibration pulsing completed
- Antenna positions marked
- Magnetic field mapped at low current
- Target installed in place
- Moved DAQ computers to counting house
- DAQ debugging continues
- Building cleaned
- Ready for the beam
21 Jan 2014 - Status as of Jan 21General:
- Water for the magnet cooling is hooked up. Few minor leaks found.
- Hang tests for UHF and VHF antenna continue
21 Jan 2014 - Photo galleries are up
The photo and data galleries are now up and running. We will fill them as we go.
20 Jan 2014 - Website is up
The website for the T-510 experiment is up. It has been created using Jekyll.
20 Jan 2014 - Status as of Jan 20General:
- DAQ system multiple waveform readout is fixed.
- First pulser test data is available.
- Another hang test for the UHF antenna tower is conducted on the crane.
- Decided on the scheme for the tilt angle management.
- Connections for the magnet water cooling system are all in place.
- Half of the power cables are connected.
- A grid for the magnetic field mapping is drawn on the target support.
- The first hang test for the VHF antenna mounts is done on the crane.
16 Jan 2014 - Status as of Jan 16General:
- Steel plates are installed, ¼” higher than designed (not a concern)
- Coils are placed into position (not attached to the plates yet)
- Some water hoses are connected
- No main pluming yet
- No power yet
- DAQ system has been set up in ESA
- Scopes are controlled and read out from the counting house
- Initial problems with a long Ethernet cable seems to be resolved
- Read out scripts are completed
- Work in progress on data transfer to a remote server and online plots generation
- First hang test
- Work started on the VHF antenna support.
- Power supplies refurbished
- Water valves replaced
- Leaks fixed (almost)
- EMI survey in ESA building is completed
- Pulser calibration system is assembled and tested
- Antenna position survey plan is developed
- T-510 run plan is submitted